OpenStudy (anonymous):

Medal Add. 8 7/10 + 2 7/15

OpenStudy (diamondboy):

do u know how to get common denominators?

OpenStudy (anonymous):

ya you add the 10 by 15 = 25

OpenStudy (alexandervonhumboldt2):

8 7/10+ 2 7/15=10+ 21/30 + 14/30=11 1/2

OpenStudy (diamondboy):

He is right @AlexandervonHumboldt2

OpenStudy (diamondboy):

did u understand?

OpenStudy (anonymous):

but there did you guys get the 10 + 21/30

OpenStudy (anonymous):

and how did you guys get 30 from

OpenStudy (alexandervonhumboldt2):

you are wrong @harz360

OpenStudy (anonymous):

can you explain where you got the 30 from

OpenStudy (diamondboy):

do u have a calculator?

OpenStudy (anonymous):

yes